OpenStudy (anonymous):

evaluate the integral x^2+X+1 everything divided by x dx from 1 to e

OpenStudy (angela210793):

\[\int\limits_{1}^{e}(x^2+x+1)dx=x^3/3+x^2/2+x\]

OpenStudy (anonymous):

u forgot that everything is divided by x

OpenStudy (angela210793):

oh..sorry

OpenStudy (anonymous):

it should be x^2 +z+1/x

OpenStudy (anonymous):

x^2+1+1/x

OpenStudy (angela210793):

if everything is divided with x then we'd have integral S(x+1+1/x)dx from 1 to e =x^2/2+x+lnx then replace x with e and then with 1 and (e^2/2+e+1)-(1/2+1+0)=....

OpenStudy (anonymous):

and ur answer will be?

OpenStudy (angela210793):

=(e^2-2e-1)/2

OpenStudy (anonymous):

i dont get how the answer give u that what happened with the 1/2+1... so lne +1?

OpenStudy (anonymous):

lne = 1?

OpenStudy (angela210793):

lne=\[\log_{e} e\]

OpenStudy (anonymous):

did you find the 'anti derivative"?

OpenStudy (anonymous):

oh yes \[F(x)=\frac{x^2}{2}+x+\ln(x)\] \[F(e)=\frac{e^2}{2}+e+1\] \[F(1)=\frac{1}{2}+1\] subtract to get \[\frac{e^2}{2}+e-\frac{1}{2}\]